How to calculate pp&E as percentage?

Current Current assets as a Nike, percentage assets as a Tesla, percentage 5/31/15, in millions of total assets 9/30/18 in millions of total assets Current assets PP&E as percentage of PP&E as percentage of Property, plant, and equipment, net Total long-term assets Total Assets 1600 IS926+5624 IS976
Answer #1

in the given table,

PP&E is provided which is a long term asset. I assumed that PP&E is included in long term assets .

Therefore total assets [ NIKE] = current assets + Long term assets = 15,976 + 5,624 = $ 21,600.

Note: If you did not included PP&E balance in long term assets , then total assets is = 15,976 + 3,011 +5,624 = $24,611

Same logic applies for Tesla also.

Now

Current assets as percentage = current assets / total assets x 100 = 15,976 /21,600 x 100 = 73.96% or 74%

PP&E as a percenrage[assuming included in long termassets] = PP&E / total assets x 100 = 3,011/21,600 x 100 = 13.94% or 14%.

If not included in long term assets balance = 3,011/24,611 x 100 = 12.23%.

The above calculations are based on the information provided in the table..